Suppose there are no imports, taxes or other leakages in the economy. If the Marginal Propensity to Consume is .80 and the government increases spending by $40 billion, by how much would output increase in the economy?

Select one:

a. $40 billion

b. $48 billion

c. $80 billion

d. $120 billion

e. $200 billion.
Answer #1

Marginal propensity to consume, MPC = 0.80

Government increases spending by $ 40 billion

MPS-1-MPC

MPS 1-08

large MPS = {0.2}

Multiplier = 1/MPS = 1/0.2 = 5

Here, the government has increased the spending thus there will be increase in the economy

Increase = Multiplier * Spending = 5*40 = $ 200 billion.

Please contact if having any query thank you.